# Primordial black holes: constraints, potential evidence and prospects

### v1

@@ II.1 Critical collapse
The critical collapse scenario relies on the presence of large super-horizon curvature perturbations, which can naturally arise during any inflationary epoch. Producing an appreciable PBH abundance requires a substantial enhancement of the curvature perturbation power spectrum, namely $\mathcal{P}_{\zeta}(k)\sim\mathcal{O}(10^{-3}\!-\!10^{-2})$, on the scales relevant for PBH formation. While the amplitude and shape of the scalar power spectrum are tightly constrained on the largest cosmological scales by direct CMB observations [28, 29], limits on CMB spectral distortions [30, 31, 32, 33, 34] and Lyman-$\alpha$ forest observations [35], much smaller scales remain essentially unconstrained. Consequently, many primordial mechanisms have been proposed that generate a significant enhancement of the scalar spectrum at these scales (e.g., [36, 37, 38, 39, 40, 41, 42, 43, 44]). Most models capable of generating such an enhancement lead to a power spectrum that can be approximated by one of the following representative ansätze:
@@ II.2.1 Phase transitions
Cosmological first-order phase transitions can lead to formation of PBHs in bubble collisions [120, 121] or from collapsing false vacuum regions [122, 123, 124, 125]. Collapsing false vacuum regions can also arise from vacuum bubbles nucleated during inflation [126, 127, 128]. While the PBH mass function arising from the collapse of false-vacuum regions during a first-order phase transition is typically very narrow [125], the collapse of vacuum bubbles nucleated during inflation can yield a much broader distribution. In this case, the spectrum takes the form of a cut piecewise power-law,
@@
Recent attention has focused on slow and strongly supercooled first-order phase transitions, including a period of thermal inflation during which the Universe remains trapped in a metastable vacuum until well below the critical temperature [129, 130, 131, 132, 117, 133, 134]. In such transitions, the formation of only a few large bubbles per Hubble volume gives rise to considerable stochastic density fluctuations and the PBH formation is based on critical collapse. A subsequent work [134] found that, once the coupling between density perturbations and metric perturbations is properly accounted for, PBH formation is highly suppressed. However, accounting for the gradient energy of the bubble walls may change this result [135].
Large density perturbations may also emerge in strongly supercooled transitions that are not of first order [136]. Including both thermal and quantum fluctuations, Ref. [137] found that the induced curvature perturbations are, however, insufficient to source PBHs. Furthermore, to enhance PBH formation from collapsing false vacuum regions, it has been proposed that particles that are heavy in the true vacuum become trapped inside the false vacuum regions, where they are subsequently compressed into PBHs during collapse [138, 139]. However, once the backreaction of these particles on the bubble wall dynamics is taken into account, PBH formation through this mechanism is strongly suppressed [140].
@@ III.1 Dark matter
The semiclassical picture of Hawking radiation may break down as the Hawking temperature approaches the Planck scale. The consequent evolution of the BHs is uncertain, as we do not have a complete quantum theory of gravity. It has been suggested that the evaporation ends, leaving a stable Planck mass relic [180]. Furthermore, it has been suggested that the evaporation may slow down or even stop much earlier due to the memory burden effect, based on black hole information theory [181, 182].
@@ IV.1.1 Constraints
PBH evaporation constraints can be altered in models with large extra dimensions, in which the Planck mass can be substantially lowered [219, 220]. The evaporation bounds can be further modified by the memory burden effect [182] or if the collapse ends up in some exotic ultracompact object rather than a BH [221].
@@ IV.3.1 Constraints
The LIGO-Virgo-KAGRA (LVK) collaboration has detected hundreds of binary BH mergers to date [265, 266, 267, 268]. However, it has been argued that the observed merger rate is inconsistent with scenarios in which stellar-mass PBHs comprise even $0.1\%$ of the DM. Therefore, the LVK observations place a strong constraint on PBHs [269, 270, 271, 272, 273, 64, 89]. The constraint shown in Fig. 4 is based on Ref. [89], which incorporates data up to the third LVK observing run. In addition to measurements of the merger rate, unresolved PBH binaries will also induce a stochastic GW background, which implies $f_{\rm PBH}\lesssim\mathcal{O}(10^{-}2)$ for stellar-mass PBHs [272, 274]. However, this is currently weaker than the constraint obtained from resolvable PBH binary mergers.
@@ IV.4.2 Potential evidence
The main potential dynamical evidence for PBHs is the triggering of explosions of stars. In particular, some recently observed supernovæ, the so-called “calcium-rich transients”, do not trace the stellar density but are located off-centre from their host galaxies and appear to originate from white dwarfs with masses of around $0.6M_{\odot}$. Ref. [234] argues that these events could have been triggered by collisions with PBHs in the mass range $10^{-12}-10^{-10}M_{\odot}$ with $10^{-3}<f_{\rm PBH}<0.1$. The wider mass range of $10^{-14}-10^{-8}M_{\odot}$ has been invoked to explain how some $r$-process elements (i.e. those generated by fast nuclear reactions) can be produced by the interaction of PBHs with neutron stars [300]. Furthermore, Ref. [301] argues that collisions of neutron stars with PBHs of $\sim 10^{-10}M_{\odot}$ may explain the millisecond durations and large luminosities of fast radio bursts.
@@ VII.1 Scalar-induced gravitational waves
In the critical-collapse scenario, the same curvature perturbations that lead to PBH formation also generate scalar-induced gravitational waves (SIGWs). More generally, SIGWs are produced whenever curvature perturbations re-enter the horizon, even if their amplitude falls below the threshold required for PBH formation. The shape of the resulting GW spectrum is determined by the properties and statistics of the underlying scalar perturbations [356, 357, 358, 359, 360, 361, 362, 363, 364, 365]. On CMB scales, the scalar spectrum is tightly constrained, and SIGWs have undetectably small amplitudes. However, PBH production requires a boost in the power spectrum at smaller scales, and this would inevitably enhance the SIGW signal. In this section, we focus on the interplay between SIGWs and PBHs, exploiting their common origin and relating current and future GW data to PBH production (see also [366, 367, 368, 369, 370, 371]).
@@ VII.1.2 Imprints of non-Gaussianity
A noteworthy aspect is that SIGWs are not only sensitive to the expansion history of the Universe, but also to primordial NG, as is evident from Eq. (LABEL:eq::SIGW_spectrum), where the tensor spectrum is dependent on the trispectrum of scalar perturbations [380, 381, 382, 383, 384, 385, 386, 387, 388, 389]. We argued in Sec. II.1.2 that NG can have an enormous impact on the PBH abundances. By contrast, the impact of NG on the SIGW is not fully understood [112]. Regardless of how one includes NG in the computation, it is expected to modify the GW spectrum, not only enhancing or suppressing it but also leaving specific spectral features. Explicitly, this can be understood by decomposing the trispectrum into a connected and a disconnected component $\mathcal{T}(\textbf{k}_{1},\textbf{k}_{2},{\textbf{k}_{3}},{\textbf{k}_{4}})=\mathcal{T}(\textbf{k}_{1},\textbf{k}_{2},{\textbf{k}_{3}},{\textbf{k}_{4}})|_{c}+\mathcal{T}(\textbf{k}_{1},\textbf{k}_{2},{\textbf{k}_{3}},{\textbf{k}_{4}})|_{d}$. In the absence of NG, the connected contribution vanishes and only the disconnected contribution survives, resulting in a term proportional to products of curvature power spectra $\mathcal{P}_{\zeta}(k_{1})\mathcal{P}_{\zeta}(k_{2})$. A simple way to model primordial NG is to adopt the local ansatz of Eq. (16). As in the PBH case, the use of this ansatz has many limitations [112], but it nevertheless reveals the effects of NG on the GW spectrum [390, 377, 391, 378]. We stress that this affects the GW spectrum quite differently from the PBH abundance. In the latter case, it substantially changes the tail of the distribution, inducing a variation in the abundance of many orders of magnitude. For SIGWs, the main effect is the presence of additional features in the GW spectrum with a relatively small variation in the amplitude.
@@ VII.3 Lensing of gravitational waves
In the solar-mass range, optical ML searches lose sensitivity because the potential events become longer than the monitoring time. However, for GW signals, the coherent nature of the waveform allows the detection of lensing effects even when the lens motion is negligible. In this regime, lensing produces a frequency-dependent magnification that imprints a characteristic frequency dependence on the signal’s amplitude [405]. In fact, the event GW231123 [406] has been recently claimed to be a potential candidate for the first lensed GW event [407].

### v2

@@ II.1 Critical collapse
The critical collapse scenario relies on the presence of large super-horizon curvature perturbations, which can naturally arise during any inflationary epoch. Producing an appreciable PBH abundance requires a substantial enhancement of the curvature perturbation power spectrum, namely $\mathcal{P}_{\zeta}(k)\sim\mathcal{O}(10^{-3}\!-\!10^{-2})$, on the scales relevant for PBH formation. While the amplitude and shape of the scalar power spectrum are tightly constrained on the largest cosmological scales by direct CMB observations [28, 29], limits on CMB spectral distortions [30, 31, 32, 33, 34] and Lyman-$\alpha$ forest observations [35], much smaller scales remain essentially unconstrained. Consequently, many primordial mechanisms have been proposed that generate a significant enhancement of the scalar spectrum at these scales (e.g., [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]). Most models capable of generating such an enhancement lead to a power spectrum that can be approximated by one of the following representative ansätze:
@@ II.2.1 Phase transitions
Cosmological first-order phase transitions can lead to formation of PBHs in bubble collisions [122, 123] or from collapsing false vacuum regions [124, 125, 126, 127]. Collapsing false vacuum regions can also arise from vacuum bubbles nucleated during inflation [128, 129, 130, 131]. While the PBH mass function arising from the collapse of false-vacuum regions during a first-order phase transition is typically very narrow [127], the collapse of vacuum bubbles nucleated during inflation can yield a much broader distribution. In this case, the spectrum can take the form of a cut piecewise power-law,
@@
Recent attention has focused on slow and strongly supercooled first-order phase transitions, including a period of thermal inflation during which the Universe remains trapped in a metastable vacuum until well below the critical temperature [132, 133, 134, 135, 119, 136, 137, 138]. In such transitions, the formation of only a few large bubbles per Hubble volume gives rise to considerable stochastic density fluctuations and the PBH formation is based on critical collapse. A subsequent work [137] found that, once the coupling between density perturbations and metric perturbations is properly accounted for, PBH formation is highly suppressed. However, accounting for the gradient energy of the bubble walls may change this result [139].
Large density perturbations may also emerge in strongly supercooled transitions that are not of first order [140]. Including both thermal and quantum fluctuations, Ref. [141] found that the induced curvature perturbations are, however, insufficient to source PBHs. Furthermore, to enhance PBH formation from collapsing false vacuum regions, it has been proposed that particles that are heavy in the true vacuum become trapped inside the false vacuum regions, where they are subsequently compressed into PBHs during collapse [142, 143]. However, once the backreaction of these particles on the bubble wall dynamics is taken into account, PBH formation through this mechanism is strongly suppressed [144] unless the trapped particles possess sufficiently strong attractive self-interactions [145].
@@ III.1 Dark matter
The semiclassical picture of Hawking radiation may break down as the Hawking temperature approaches the Planck scale. The consequent evolution of the BHs is uncertain, as we do not have a complete quantum theory of gravity. It has been suggested that the evaporation ends, leaving a stable Planck mass relic [185]. Furthermore, it has been suggested that the evaporation may slow down or even stop much earlier due to the memory burden effect, based on black hole information theory [186, 187, 188, 189, 190].
@@ IV.1.1 Constraints
PBH evaporation constraints can be altered in models with large extra dimensions, in which the Planck mass can be substantially lowered [227, 228]. The evaporation bounds can be further modified by the memory burden effect [187, 190] or if the collapse ends up in some exotic ultracompact object rather than a BH [229].
@@ IV.3.1 Constraints
The LIGO-Virgo-KAGRA (LVK) collaboration has detected hundreds of binary BH mergers to date [273, 274, 275, 276]. However, it has been argued that the observed merger rate is inconsistent with scenarios in which stellar-mass PBHs comprise even $0.1\%$ of the DM. Therefore, the LVK observations place a strong constraint on PBHs [277, 278, 279, 280, 281, 66, 91, 282]. The constraint shown in Fig. 4 is based on Ref. [91], which incorporates data up to the third LVK observing run. In addition to measurements of the merger rate, unresolved PBH binaries will also induce a stochastic GW background, which implies $f_{\rm PBH}\lesssim\mathcal{O}(10^{-}2)$ for stellar-mass PBHs [280, 283]. However, this is currently weaker than the constraint obtained from resolvable PBH binary mergers.
@@ IV.4.2 Potential evidence
The main potential dynamical evidence for PBHs is the triggering of explosions of stars. In particular, some recently observed supernovæ, the so-called “calcium-rich transients”, do not trace the stellar density but are located off-centre from their host galaxies and appear to originate from white dwarfs with masses of around $0.6M_{\odot}$. Ref. [242] argues that these events could have been triggered by collisions with PBHs in the mass range $10^{-12}-10^{-10}M_{\odot}$ with $10^{-3}<f_{\rm PBH}<0.1$. The wider mass range of $10^{-14}-10^{-8}M_{\odot}$ has been invoked to explain how some $r$-process elements (i.e. those generated by fast nuclear reactions) can be produced by the interaction of PBHs with neutron stars and this may also explain the large luminosities and millisecond durations of fast radio bursts (FRBs) [309, 310]. Estimates of the rates of such collisions indicate that PBHs can account for the observed FRBs only if the PBH density profile is sufficiently spiky in galactic centers [311, 312].
@@ VII.1 Scalar-induced gravitational waves
In the critical-collapse scenario, the same curvature perturbations that lead to PBH formation also generate scalar-induced gravitational waves (SIGWs). More generally, SIGWs are produced whenever curvature perturbations re-enter the horizon, even if their amplitude falls below the threshold required for PBH formation. The shape of the resulting GW spectrum is determined by the properties and statistics of the underlying scalar perturbations [367, 368, 369, 370, 371, 372, 373, 374, 375, 376]. On CMB scales, the scalar spectrum is tightly constrained, and SIGWs have undetectably small amplitudes. However, PBH production requires a boost in the power spectrum at smaller scales, and this would inevitably enhance the SIGW signal. In this section, we focus on the interplay between SIGWs and PBHs, exploiting their common origin and relating current and future GW data to PBH production (see also [377, 378, 379, 380, 381, 382, 383]).
@@ VII.1.2 Imprints of non-Gaussianity
A noteworthy aspect is that SIGWs are not only sensitive to the expansion history of the Universe, but also to primordial NG, as is evident from Eq. (LABEL:eq::SIGW_spectrum), where the tensor spectrum is dependent on the trispectrum of scalar perturbations [392, 393, 394, 395, 396, 397, 398, 399, 400, 401]. We argued in Sec. II.1.2 that NG can have an enormous impact on the PBH abundances. By contrast, the impact of NG on the SIGW is not fully understood [114]. Regardless of how one includes NG in the computation, it is expected to modify the GW spectrum, not only enhancing or suppressing it but also leaving specific spectral features. Explicitly, this can be understood by decomposing the trispectrum into a connected and a disconnected component $\mathcal{T}(\textbf{k}_{1},\textbf{k}_{2},{\textbf{k}_{3}},{\textbf{k}_{4}})=\mathcal{T}(\textbf{k}_{1},\textbf{k}_{2},{\textbf{k}_{3}},{\textbf{k}_{4}})|_{c}+\mathcal{T}(\textbf{k}_{1},\textbf{k}_{2},{\textbf{k}_{3}},{\textbf{k}_{4}})|_{d}$. In the absence of NG, the connected contribution vanishes and only the disconnected contribution survives, resulting in a term proportional to products of curvature power spectra $\mathcal{P}_{\zeta}(k_{1})\mathcal{P}_{\zeta}(k_{2})$. A simple way to model primordial NG is to adopt the local ansatz of Eq. (16). As in the PBH case, the use of this ansatz has many limitations [114], but it nevertheless reveals the effects of NG on the GW spectrum [402, 403, 389, 404, 390]. We stress that this affects the GW spectrum quite differently from the PBH abundance. In the latter case, it substantially changes the tail of the distribution, inducing a variation in the abundance of many orders of magnitude. For SIGWs, the main effect is the presence of additional features in the GW spectrum with a relatively small variation in the amplitude.
@@ VII.3 Lensing of gravitational waves
In the solar-mass range, optical ML searches lose sensitivity because the potential events become longer than the monitoring time. However, for GW signals, the coherent nature of the waveform allows the detection of lensing effects even when the lens motion is negligible. In this regime, lensing produces a frequency-dependent magnification that imprints a characteristic frequency dependence on the signal’s amplitude [418]. In fact, the event GW231123 [419] has been recently claimed to be a potential candidate for the first lensed GW event [420, 421, 422, 423].